# RUNHETC-2025-36 JT de Sitter Gravity as a Model of Coleman-de Luccia Tunneling

### v2

@@ 1 Introduction
The dilaton field blows up at $\chi=\pm\infty$ and the solution is thus eliminated by any reasonable boundary conditions on the other solutions, such as those that follow naturally from analytic continuation. Since the Poincaré solution is just a half-coordinate patch of the global solution, related by a coordinate transformation, the global solution is also ruled out by the boundary conditions. As a consequence, all of the Lorentzian solutions are invariant under translations in $\chi$ and we are free to view this as a gauge symmetry of the problem even though the range of the variable is non-compact. This will no longer be true if we couple the model to systems with true field theoretic degrees of freedom. Below we describe formalisms for treating both the possibility of translation gauge symmetry and one in which there might be excitations that are localized in $\chi$.
@@ 2 The Instanton and Its Continuations
Positivity of entropy implies $S_{0}\pm S_{1}\geq 0$, where 00 indicates a singularity. We assume that, it is at the left pole $\theta=0$ that the static patch transitions into the expanding dS cosmology, and at the right pole $\theta=\pi$ that it transitions into a Big Crunch.
@@
Time dependent Hamiltonians in quantum gravity are the analog of half sided modular inclusions in quantum field theory. If one believes in the idea of a causal diamond as a subsystem and that the covariant entropy bound has any precise quantum mechanical meaning, then operator algebras associated with diamond subsystems cannot be Type III. This means that the only way that diamonds can truly be subsystems [^2] In a lattice field theory with a time independent Hamiltonian, a finite region of space is a subsystem only for an infinitesimal period of time. Any finite time translation operator connects it to regions that are arbitrarily far away, albeit with exponentially small amplitude. is if the Hamiltonian is time dependent. If we think of diamond Hilbert spaces as growing in dimension with the area of the diamond’s holographic screen, then the time dependent mapping is a unitary embedding of a small diamond Hilbert space into that of a larger diamond, the proper time from some common origin of whose future tip, along some particular geodesic, is larger by one Planck unit than the smaller diamond (see Fig. 1 in Appendix). This embedding must be supplemented by a unitary transformation in the commutant of the larger diamond operator algebra, to give a unitary on the full Hilbert space of the model.
@@
Now imagine the system in the metastable state represented by the static patch solution. Notice that unlike the conventional static coordinates on dS space, this coordinate patch has two horizons, of differing entropy. This is analogous to the Schwarzschild-de Sitter black hole in $4$ or more dimensions. That solution has perturbative instabilities indicated by the higher temperature of the black hole horizon. It emits gravitons and disappears back to the stable equilibrium of empty dS space. In JTdS there are no particle excitations to emit and the low entropy horizon decays by tunneling. The system it decays into has a time dependent density matrix. At early times its density matrix is a tensor product of a large number of copies of a density matrix of very low entropy [^3] The full Hilbert space of JTdS gravity has infinite dimension, as indicated by the asymptotic value of $S(\tau)$ in the Milne solution. When we refer to tensor products we are always imagining the causal diamond Hilbert space at some large time $T$, where $T/L_{P}$ is taken to infinity at the end of the calculation. We assume that number theoretic issues having to do with when we can factorize the dimension of the full Hilbert space into products of the dimensions of finite diamond Hilbert spaces, become irrelevant in the limit.
@@ 3 Quantum Mechanical Models
where we’ve used an obvious notation for the un-normalized weight matrices. We now assume that the modular Hamiltonians of the two systems belong to the large N universality class described above, which are well approximated by operators quadratic in fermions. In this case the traces are products of a large number of factors. The low entropy constraint on the static patch simply means that most of those factors in the numerator are simply $=1$, while the corresponding denominator factors are $1+e^{-\lambda_{a}}$. Therefore, the ratio is

### v3

@@ 1 Introduction
The dilaton field blows up at $\chi=\pm\infty$ and the solution is thus eliminated by any reasonable boundary conditions on the other solutions, such as those that follow naturally from analytic continuation. Since the Poincaré solution is just a half-coordinate patch of the global solution, related by a coordinate transformation, the global solution is also ruled out by the boundary conditions. As a consequence, all of the Lorentzian solutions are invariant under translations in $\chi$ and we are free to view this as a gauge symmetry of the problem even though the range of the variable is non-compact. This will no longer be true if we couple the model to systems with true field theoretic degrees of freedom. In the following sections, we describe formalisms for treating both the possibility of translation gauge symmetry and one in which there might be excitations that are localized in $\chi$.
Throughout this paper we will be interpreting the $S$ function as describing the entropy in some causal diamond. This is in accord with what we call the Covariant Entropy Principle (CEP) in higher dimensions, and the fact that $S$ arises from the volume of a compact d2d-2 fold when a two dimensional theory is obtained from a $d$ dimensional one by compactification. The CEP states that the entropy in a causal diamond is $\frac{A_{\diamond}}{4G_{N}}$, where $A_{\diamond}$ is the maximal d2d-2 volume on the diamond boundary. This principle can be viewed as lying behind Jacobson’s [24] derivation of Einstein’s equations from hydrodynamics. For black holes it has been “derived” by Carlip [25] and Solodukhin [26], and [27] extended that derivation to arbitrary causal diamonds. For causal diamonds in maximally symmetric space-times the CEP has been derived by Euclidean path integral calculations [28]. At any rate, it is the basic assumption behind our work.
@@ 2 The Instanton and Its Continuations
Positivity of entropy implies $S_{0}\pm S_{1}\geq 0$, where 0 indicates a singularity. We assume that, it is at the left pole $\theta=0$ that the static patch transitions into the expanding dS cosmology, and at the right pole $\theta=\pi$ that it transitions into a Big Crunch.
@@
Time dependent Hamiltonians in quantum gravity are the analog of half sided modular inclusions in quantum field theory. If one believes in the idea of a causal diamond as a subsystem and that the covariant entropy principle has any precise quantum mechanical meaning, then operator algebras associated with diamond subsystems cannot be Type III. There are indications [36, 37] that systems with time independent Hamiltonians can only have sharp light cones if the spectrum of the Hamiltonian admits infinite energy states. This means that the only way that diamonds with finite dimensional Hilbert spaces can truly be subsystems [^2] Take a lattice field theory with a time independent Hamiltonian as an example, a finite region of space in such a theory is a subsystem only for an infinitesimal period of time. Any finite time translation operator connects it to regions that are arbitrarily far away, albeit with exponentially small amplitude. is if the Hamiltonian is time dependent. We note that the time-band sub-algebras of Liu and Leutheusser [38] exist only in the strict $N=\infty$ limit, and are Type III.
With time dependent unitary evolution it’s easy to construct exactly causal dynamics [39, 40, 41, 42, 43]. These flows can be thought of as quantum gravity generalizations of the one sided modular inclusions of algebraic quantum field theory. If we think of diamond Hilbert spaces as growing in dimension with the area of the diamond’s holographic screen, then the time dependent mapping is a unitary embedding of a small diamond Hilbert space into that of a larger diamond, the proper time from some common origin of whose future tip, along some particular geodesic, is larger by one Planck unit than the smaller diamond (see Fig. 1 in Appendix). This embedding must be supplemented by a unitary transformation in the commutant of the larger diamond operator algebra, to give a unitary on the full Hilbert space of the model.
This sort of time evolution has been used extensively in holographic space-time models for almost a quarter of a century [39, 40, 41, 42, 43]. At each time, one factors the Hilbert space into ${\cal H}_{in}\times{\cal H}_{out}$. The time evolution operator is written as
$U(t)=U_{in}(t)\otimes U_{out}(t).$ (11)
$U_{in}(t)$ is the unitary embedding between consecutive nested diamond Hilbert spaces described above (the analog of one sided modular inclusion). The authors of [39, 40, 41, 42, 43] conjectured that $U_{out}(t)$ is determined by the Quantum Principle of Relativity: One constructs independent time evolutions along all the geodesics in some background metric, and requires that the dynamics be such that one can consistently assign the a density matrix with the same entanglement spectrum to the maximal causal diamond in the overlap between any two diamonds along an arbitrary pair of geodesics. Further details can be found in [39, 40, 41, 42, 43].
@@
Now imagine the system in the metastable state represented by the static patch solution. Notice that unlike the conventional static coordinates on dS space, this coordinate patch has two horizons, of differing entropy. This is analogous to the Schwarzschild-de Sitter black hole in $4$ or more dimensions. That solution has perturbative instabilities indicated by the higher temperature of the black hole horizon. It emits gravitons and disappears back to the stable equilibrium of empty dS space. In JTdS there are no particle excitations to emit and the low entropy horizon decays by tunneling. The system it decays into has a time dependent density matrix. At early times its density matrix is a tensor product of a large number of copies of a density matrix of very low entropy [^3] The full Hilbert space of JTdS gravity has infinite dimension, as indicated by the asymptotic value of $S(\tau)$ in the Milne solution. When we refer to tensor products we are always imagining the causal diamond Hilbert space at some large time $T$, where $T/L_{P}$ is taken to infinity at the end of the calculation. We assume that number theoretic issues having to do with when we can factorize the dimension of the full Hilbert space into products of the dimensions of finite diamond Hilbert spaces, become irrelevant in the limit.. This is the usual assumption that there are no correlations between causally disconnected regions of the universe built into the initial conditions of cosmology. Without such an assumption the issue of “the Horizon Problem” would never arise.
@@ 3 Quantum Mechanical Models
where we’ve used an obvious notation for the un-normalized weight matrices. We now assume that the modular Hamiltonians of the two systems belong to the large N universality class described above, which are well approximated by operators quadratic in fermions. In this case the traces are products of a large number of factors. The low entropy constraint on the static patch simply means that most of those factors in the numerator are simply =1=1, while the corresponding denominator factors are $1+e^{-\lambda_{a}}$. Therefore, the ratio is